N IS A NATURAL NUMBER,1 is written after the five digit number to make it as asix digit number,which is thrice the same number if one is written after it,find the number

Grade:10

1 Answers

Shobhit Varshney IIT Roorkee
8 years ago
Dear student,

10N + 1 = 3(105 + N)
10N + 1 = 3*105 + 3N
7N = 299999
N = 42857
….….….….….….….….….….….….….….….….….…....
Thanks